# Conditional Quantile Estimation for Uncertain Watch Time in Short-Video Recommendation

### v1

@@ 3.3.1. Conditional Expectation
The most intuitive strategy is recovering the mean estimation through the conditional expectation. One challenge of this strategy is that we do not have the output value for $\tau\in(\tau_{i},\tau_{i+1})$ between any two consecutive quantiles. To circumvent this lack of information, we can use the interpolation method to approximate the conditional distribution. In practice, we adopt a linear interpolation between consecutive quantiles so the expected watch-time between $\tau_{i}$ and $\tau_{i+1}$ becomes $(t_{\tau_{i}}+t_{\tau_{i+1}})/2(N+1)$. For the two endpoints, we assume $t_{0}=t_{\tau_{1}}$ and $t_{1}=t_{\tau_{N}}$. Then, we can approximate the overall watch-time expectation as:
(3) $\mathbb{E}[W|\mathbf{x}]$$\approx$ $\frac{1}{2(N+1)}[(t_{\tau_{1}}+t_{\tau_{1}})+(t_{\tau_{1}}+t_{\tau_{2}})++(t_{\tau_{2}}+t_{\tau_{3}})...$
$+(t_{\tau_{N-2}}+t_{\tau_{N-1}})+(t_{\tau_{N-1}}+t_{\tau_{N}})+(t_{\tau_{N}}+t_{\tau_{N}})]$
$=$ $\frac{1}{N+1}\sum_{i=1}^{N}t_{\tau_{i}}+\frac{t_{\tau_{1}}+t_{\tau_{N}}}{2(N+1)}.$
@@
Theoretically, this expectation gives the most accurate prediction in general and it would achieve the optimal prediction when $N\rightarrow\infty$. Empirically, we will verify its superiority with our experimental analysis in section 4. Yet, we remind readers that this strategy may not be well-suited for scenarios where users are not tolerant of bad recommendations or the recommender system requires dynamic controls. Fortunately, the proposed CQE is a flexible framework that provides sufficient freedom on how a strategy would combine different quantile outputs.
3.3.2. Conservative Estimations
In environments with implicit uncertainty, especially online platforms where users are sensitive to bad recommendations, conservative estimation stands as a pivotal strategy. For example, a video with high average watch-time may actually being polarized towards 0s and 1s in its watch-time distribution, so the expectation is no longer representative in this case. Adopting a conservative approach in video recommendations means that we should favor a quantile that reveals the underestimated watch-time scenarios, in contrast with an optimistic approach that selects overestimated watch-time. Consequently, the inferred watch-time is likely to be exceeded by the actual user behavior and the recommendation ensures that the lower bounded watch-time is still satisfactory.
Formally, for a user-video pair, we focus on the lower quantiles of the expectation Eq.(3) and select one of the quantile $\tau_{\text{low}}$ as the inference. In practice, $\tau_{\text{low}}$ may be set to reflect our degree of risk aversion. For instance, a $\tau_{\text{low}}$ of 0.25 could be used if we wish to ensure that the actual watch-time exceeds the predicted time at least 75% of the time. It is more cautious than using the median and enables a recommendation strategy that systematically avoids user disappointment due to overly optimistic estimates.
3.3.3. Dynamic Quantile Combination
Intuitively, user engagement is a multifaceted behavior influenced by the interplay of various user-specific and content-specific factors, which results in the different propensity towards longer or shorter watch-time. Following this ever-changing nature of user interest and platform content, we also designed a dynamic quantile combination (DQC) strategy that allows the system to control the combination between the higher and lower quantile. Specifically, we use a blending parameter $k$ that is carefully designed to encode the live feedback and activity level of users, as well as the novelty and diversity of the recommended content. Then the combined score $s_{\text{dyn}}$ is computed as a weighted average of lower and upper quantile predictions:
(4) $s_{\text{dyn}}=t_{\tau_{\text{low}}}\cdot k+t_{\tau_{\text{high}}}\cdot(1-k).$
where $t_{\tau_{\text{low}}}$ represent a conservative quantile prediction and $t_{\tau_{\text{high}}}$ corresponds to a more optimistic one. The blending parameter $k$ is thus a tuning knob that focuses the recommendation’s strategy between these two bounds.
@@ 4. Experiments and Results
We conduct extensive experiments in both offline and online environments to demonstrate the effectiveness of CQE. Three research questions are investigated in the experiments:
@@
  • How do Conditional Expectation (CDE) of Conditional Quantile Estimation (CQE) compare with the state-of-the-art methods in terms of predicting user interest and watch time with regard to recommendation accuracy?
  • What is the impact of the number of quantiles on the performance of CQE-CED?
  • How do the strategies Conditional Expectation(CDE), Conservative Estimation(CSE), and Dynamic Quantile Combination(DQC) perform when implemented on online platforms?
4.1. Offline Experiments
Our offline experiments were designed in line with two existing research settings. The first focused directly on watch-time prediction as proposed by TPM (Lin et al., 2023). The second aimed at predicting user interest through duration bias adjustment as $\text{D}^{2}$o(Zhao et al., 2023).
4.1.1. Watch-Time Prediction.
In this problem scope, our primary objective is to predict the watch-time. We employed the same training methodology as referred to in the TPM code.
Datasets. Following TPM, we used two public datasets: Kuaishou (collected from Kuaishou App [^1] https://kuairec.com/) and CIKM16 (from CIKM16 Cup  [^2] https://competitions.codalab.org/competitions/11161) for our experiments. CIKM16 aims to predict the dwell time for each session in online search results. Each item in the session is used as a single feature for input. Kuaishou dataset contains 7,176 users, 10,728 items, and 12,530,806 impressions; and CIKM16 dataset contains 310,302 sessions, and 122,991 items and the average length of each session is 3.981.
  • MAE: This metric is a typical measurement for evaluating regression accuracy. Denote the prediction as $\hat{y}$ and the true watch time as $y$,
    (5) $\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{y}_{i}-y\|$
  • XAUC: this metric evaluates if the predictions of two samples are in the same order with their true watch time. Such pairs are uniformly sampled and the percentile of samples that are correctly ordered by predictions is XAUC.
Baselines. Four state-of-the-art methods for watch time prediction were selected for comparison, including WLR(Weighted Logistic Regression) (Covington et al., 2016), D2Q(Duration-Deconfounded Quantile) (Zhan et al., 2022), OR(Ordinal Regression) (Niu et al., 2016) and TPM(Tree-based Progressive regression Model) (Lin et al., 2023). The first three methods are deterministic, while the latter introduces an element of uncertainty, offering both an estimate of the mean and the variance.
Table 1. Comparison between CDE-CQE and other baselines in user interest prediction task. Boldface means the best-performed methods in each backbone.
Backbone Method KuaiRand WeChat
GAUC nDCG@1 nDCG@3 nDCG@5 GAUC nDCG@1 nDCG@3 nDCG@5
DeepFM PCR + MSE 0.592 0.411 0.465 0.502 0.537 0.565 0.572 0.578
PCR + CE 0.635 0.435 0.486 0.521 0.553 0.584 0.586 0.591
PCR + CDE-CQE(Ours) 0.638 0.438 0.488 0.523 0.593 0.654 0.649 0.649
WTG + MSE 0.601 0.417 0.470 0.506 0.537 0.578 0.579 0.585
WTG + CE 0.647 0.440 0.491 0.526 0.545 0.565 0.569 0.576
WTG + CDE-CQE(Ours) 0.654 0.444 0.495 0.529 0.577 0.621 0.622 0.625
$\text{D}^{2}$o + MSE 0.606 0.416 0.473 0.509 0.528 0.578 0.577 0.580
$\text{D}^{2}$o + CE 0.656 0.447 0.496 0.530 0.531 0.570 0.571 0.576
$\text{D}^{2}$o + CDE-CQE(Ours) 0.662 0.452 0.500 0.533 0.557 0.635 0.626 0.625
AutoInt PCR + MSE 0.595 0.414 0.467 0.503 0.536 0.565 0.572 0.578
PCR + CE 0.639 0.437 0.488 0.522 0.554 0.579 0.583 0.588
PCR + CDE-CQE(Ours) 0.640 0.441 0.489 0.524 0.593 0.655 0.649 0.649
WTG + MSE 0.602 0.419 0.470 0.507 0.536 0.568 0.573 0.578
WTG + CE 0.648 0.441 0.492 0.526 0.546 0.563 0.570 0.576
WTG + CDE-CQE(Ours) 0.652 0.448 0.495 0.528 0.576 0.631 0.625 0.627
$\text{D}^{2}$o + MSE 0.610 0.420 0.474 0.510 0.527 0.582 0.580 0.583
$\text{D}^{2}$o + CE 0.658 0.447 0.497 0.531 0.531 0.570 0.571 0.576
$\text{D}^{2}$o + CDE-CQE(Ours) 0.661 0.455 0.500 0.534 0.557 0.635 0.626 0.624
4.1.2. User Interest Prediction.
In this context, the training label is based on the transformation of watch time and duration, whereas the testing label is the user interest indicator defined as long_view. Following $\text{D}^{2}$o (Zhao et al., 2023), in detail, it defines the user interest for a given user-video pair $(u,v)$ as:
where $d$ is video duration and $w$ is watch-time. We adopted the same training configuration as $\text{D}^{2}$o and used the classical deep recommendation model DeepFM (Guo et al., 2017) and the state-of-the-art recommendation model AutoInt (Song et al., 2019) as our backbone recommendation model.
where $\sigma$ is the sigmoid function and $r$ is the user’s interest defined by PCR, WTG or $\text{D}^{2}$o. Following $\text{D}^{2}$o, in PCR and WTG, we treat all samples with less than 5 seconds of watch time as 0 values after calculating the value of labels. This can help remove the noise in watch time.
By default, we set the number of quantiles $N$ to 100. The value of $\tau_{low}$ is selected empirically from 0.2, 0.25, and 0.3. Similarly, the value of $\tau_{high}$ is selected empirically from 0.6, 0.7, and 0.8.
Table 2. Comparison between CDE-CQE and other approaches in watch-time prediction task. Boldface means the best-performed methods.
Methods Kuaishow CIKM16
MAE XAUC MAE XAUC
WLR 6.047 0.525 0.998 0.672
D2Q 5.426 0.565 0.899 0.661
OR 5.321 0.558 0.918 0.664
TPM 4.741 0.599 0.884 0.676
CDE-CQE(Ours) 4.437 0.610 0.823 0.694
4.1.3. Experimental Results
We summarize the results as follows:
Comparison between CDE-CQE and other methods: We compared the performances of different approaches in the watch-time prediction task, and the results are listed in Table 2. Both TPM and CDE-CQE outperform other methods in both MAE and XAUC metrics, thereby highlighting the significance of incorporating uncertainty into the models. Furthermore, our approach demonstrates superior performance on both metrics compared to TPM, thereby emphasizing the advantages of employing quantile modeling techniques. Additionally, the consistent behavior between MAE and XAUC metrics also verifies the feasibility of the watch-time estimation serving as ranking metrics. As for the user interest prediction task, we compare different frameworks across backbone models (DeepFM and AutoInt) and various label designs (PCR, WTG, and $\text{D}^{2}$o) and present the results in Table 1, our proposed CDE-CQE consistently outmatches the alternatives in all cases, indicating the robustness and effectiveness of CDE-CQE. In terms of the optimization framework, CE generally performs better than MSE, indicating the correctness of including the ordinal categorical information as guidance. And CDE-CQE can improve over CE across all the designs of user interest metrics (PCR, WTG, and $\text{D}^{2}$o), which means that the proposed framework is generalizable to different label settings.
The Effect of Hyper-Parameters in CDE-CQE: To better investigate the characteristics of the proposed CQE framework, we further conduct an ablation study on the number of quantiles $N$ by varying its value from 1 to 500. Theoretically, larger $N$ generates a more accurate approximation of the truth expectation and in turn achieves better recommendation performance in general. This is attributed to the fact that more quantiles yield a distribution that closely mirrors the actual one. In the context of the watch time prediction task, as evidenced by Figure 2, the model performance improves with an increase in predicted quantiles. In contrast, for the user interest prediction task, it was observed (shown in 3) that model performance was relatively weak when the number of quantiles was under 10. Beyond 10, the results oscillated around 0.663. Interestingly, unlike the watch time prediction task, more quantiles did not necessarily lead to better results. This discrepancy suggests that there is a gap between the training objective and the user interest label as defined in the testing set. In practice, another critical issue that may hinder the employment of large $N$ is the trade-off between the marginal accuracy improvement and the linearly increased computational cost. In general, increasing $N$ would potentially improve the prediction accuracy under the Conditional Expectation strategy, but there may exists several bottlenecks that potentially limit the performance: 1) the design of the target label; 2) the data characteristics; and 3) the computational concerns.
4.2. Online Experiments
Figure 2. The performances of CDE-CQE with various number of quantiles in CIKM16 Dataset for predicting watch time.
[…]
A woman and a girl in white dresses sit in an open car.
Figure 3. The performances of CDE-CQE with various number of quantiles in KuaiRand Dataset for predicting user interest.
[…]
A woman and a girl in white dresses sit in an open car.
@@ 4.2. Online Experiments
To validate the real-world impact of our Conditional Quantile Estimation (CQE) framework, we conducted extensive online A/B tests on the KuaiShou platform, which serves over 300 million daily active users. These experiments allowed us to assess the practical effectiveness of all three CQE strategies in a live environment with a substantial user base.
@@ 4.2.1. Experiment Setup
Users were randomly assigned to control and experimental groups, with a minimum of 10% of daily user traffic allocated to the experimental group to ensure statistical significance. Each online A/B test ran for over a week, providing ample time for data collection and reliable result analysis.
The recommendation system operates in a two-stage process: candidate retrieval followed by ranking. We incorporated the CQE model into the ranking stage to predict watch time, a crucial component of the recommendation process.
@@
4.2.2. Experiment Results
Table LABEL:tab:online_results and Table LABEL:tab:user_retention summarize the performance of the CQE strategies relative to the baseline across all four metrics:
@@ 4.2.2. Experiment Results
Conservative Estimation (CSE): CSE yielded a balanced improvement across all metrics. It achieved a modest 0.008% increase in ”Average Watch Time per User” while simultaneously boosting ”Total Play Count” by 0.346%. Furthermore, CSE led to a 0.033% increase in active days and a 0.031% growth in active users. These results indicate that CSE successfully encourages users to interact with more videos, visit the platform more frequently, and maintain longer-term engagement, albeit with slightly shorter average watch times per video.
Dynamic Quantile Combination (DQC): By setting the blending parameter $k$ based on content novelty, DQC achieved improvements across both engagement and interaction metrics. It increased ”Average Watch Time per User” by 0.106% and ”Total Play Count” by 0.177%. While we don’t have specific data for active days and users for DQC, Figure 4 illustrates that it increased two core diversity metrics over the course of the experiment.
@@
  • DQC offers a middle ground, improving both depth and breadth of engagement while also enhancing content diversity.
@@
Table 3. CSE’s enhancements in active days and active users. A boldface means a statistically significant result (p-value $<$ 5%, confidence level $>$ 95%).
Active dyas Active Users
+0.033% +0.031%
Table 4. Comparison between CQE Strategies and baseline online. A boldface means a statistically significant result (p-value $<$ 5%, confidence level $>$ 95%).
Strategy Average Watch Time per User Total Play Count
CDE +0.165% -0.088%
CSE +0.008% +0.346%
DQC +0.106% +0.177%
@@
Figure 4. The diversity performances of DQC-CQE compared with baseline online.

### v4

@@ 3.4.3. Conditional Expectation
For scenarios in which our goal is to optimize the expected watch time, we employ a conditional expectation strategy. This approach estimates the mean watch time by interpolating between the predicted quantiles.
As depicted in Figure 1(d), the Conditional Expectation Estimation (CDE) strategy provides a global optimization perspective, aiming to maximize overall user engagement considering the entire watch time distribution.
@@
As shown in the left part of Figure 2, these output watch-time values exemplify the watch-time distribution. To recover the mean estimation through the conditional expectation, we face the challenge of not having output values for $\tau\in(\tau_{i},\tau_{i+1})$ between any two consecutive quantiles. To circumvent this lack of information, we use the interpolation method to approximate the conditional distribution.
We adopt a linear interpolation between consecutive quantiles, so the expected watch time between $\tau_{i}$ and $\tau_{i+1}$ becomes $(t_{\tau_{i}}+t_{\tau_{i+1}})/2(N+1)$. For the two endpoints, we assume $t_{0}=t_{\tau_{1}}$ and $t_{1}=t_{\tau_{N}}$. Then, we can approximate the overall watch-time expectation as:
(8) $\begin{split}\hat{y}_{\text{CDE}}=&\frac{1}{2(N+1)}[(t_{\tau_{1}}+t_{\tau_{1}})+(t_{\tau_{1}}+t_{\tau_{2}})+(t_{\tau_{2}}+t_{\tau_{3}})...\\ &+(t_{\tau_{N-2}}+t_{\tau_{N-1}})+(t_{\tau_{N-1}}+t_{\tau_{N}})+(t_{\tau_{N}}+t_{\tau_{N}})]\\ =&\frac{1}{N+1}\sum_{i=1}^{N}t_{\tau_{i}}+\frac{t_{\tau_{1}}+t_{\tau_{N}}}{2(N+1)}.\end{split}$
@@
Theoretically, this expectation gives the most accurate prediction in general and it would achieve the optimal prediction when $N\to\infty$. Empirically, we will verify its superiority with our experimental analysis in section 4.3. Yet, we remind readers that this strategy may not be well-suited for scenarios where users are not tolerant of bad recommendations or the recommender system requires dynamic controls.
Each of these inference strategies offers unique benefits, allowing the recommendation system to adapt to different objectives and user contexts. Using the rich information provided by our CQE model, we can make more informed and flexible recommendation decisions.
@@ 4. Experiments and Results
In this section, we present a comprehensive evaluation of our Conditional Quantile Estimation (CQE) model through both online A/B tests and offline experiments. Our experimental design aims to address several interconnected research questions:
  • RQ1: How do different CQE strategies perform in real-world scenarios?
@@
  • RQ2: How does CQE compare with state-of-the-art methods in watch time prediction and user interest modeling?
  • RQ3: What is the impact of the number of quantiles on CQE’s performance?
By exploring these questions, we seek to provide a holistic view of CQE’s capabilities, its practical impact, and its potential for generalization across different recommendation contexts.
4.1. Online Experiments (RQ1)
@@
To validate the real-world impact of our Conditional Quantile Estimation (CQE) framework, we conducted extensive online A/B tests on a major platform with hundreds of millions of daily active users, ensuring statistical significance.
@@ 4.1.1. Experiment Setup
Users were randomly assigned to either the control or experimental group, with each group receiving at least 10% of the daily user traffic. Each online A/B test ran for over a week, providing ample time for data collection and reliable result analysis.
Our CQE model utilizes an MLP network and is integrated into the ranking stage for watch time prediction. This CQE-based prediction is then combined with other existing prediction signals (such as like rate, video completion rate, etc.) to comprehensively rank the candidate videos.
@@
4.1.2. Experiment Results (RQ1)
Table 1. Conservative Estimation’s enhancements in active days and active users. A boldface means a statistically significant result (p-value $<$ 5%).
Active days Active Users
+0.033% +0.031%
Table 2. Comparison between CQE Strategies and baseline online. A boldface means a statistically significant result (p-value $<$ 5%).
Strategy Watch Time Play Count
Conservative Estimation +0.008% +0.346%
Dynamic Quantile Combination +0.106% +0.177%
Conditional Expectation +0.165% -0.088%
Table 2 and Table 1 summarize the performance of the CQE strategies relative to the baseline:
Conservative Estimation (CSE): CSE yielded a balanced improvement across all metrics. It achieved a modest 0.008% increase in ”Average Watch Time per User” while simultaneously boosting ”Total Play Count” by 0.346%. Furthermore, CSE led to a 0.033% increase in active days and a 0.031% growth in active users. Given the massive daily user base of hundreds of millions, even a tiny increase in active days/active users is considered statistically significant at the 0.02% level. These results indicate that CSE successfully encourages users to interact with more videos, visit the platform more frequently, and maintain longer-term engagement.
Dynamic Quantile Combination (DQC): By setting the blending parameter k based on content novelty, DQC achieved improvements across both engagement and interaction metrics. It increased ”Average Watch Time per User” by 0.106% and ”Total Play Count” by 0.177%. Figure 3 illustrates that it also increased two core diversity metrics over the course of the experiment.
@@ 4.1.2. Experiment Results (RQ1)
  • DQC offers a middle ground, improving both depth and breadth of engagement while enhancing content diversity.
@@
4.2. Distribution Characteristics Analysis
To provide deeper insights into the nature of watch time uncertainty and validate our approach’s applicability across diverse scenarios, we conducted additional analysis on the distribution characteristics of predicted watch times using online production data.
We defined the interquartile range (IQR) as $\frac{1}{N/2}\sum_{i=1}^{N/2}t_{\tau_{N-i+1}}-t_{\tau_{i}}$ to measure the spread of the conditional watch time distribution. Table 3 shows the relationship between IQR ranges and prediction accuracy metrics.
Table 3. Prediction performance across different IQR ranges in online environment
IQR Range UAUC MAE Sample Ratio
$[0.0,0.1]$ 0.721 0.105 0.077
$[0.1,0.2]$ 0.693 0.148 0.226
$[0.2,0.4]$ 0.631 0.197 0.543
$[0.4,0.5]$ 0.573 0.266 0.134
$[0.5,1.0]$ 0.536 0.305 0.020
As shown in Table 3, the results reveal a clear pattern: samples with narrower distribution spreads (smaller IQR) consistently achieve better prediction performance. Notably, the majority of samples (54.3%) fall within the medium IQR range $[0.2,0.4]$, while only 2% of samples exhibit very wide distributions with IQR in $[0.5,1.0]$. This validates our approach’s focus on modeling the full conditional distribution rather than a single point estimate, as the distribution spread varies significantly across different user-video pairs.
Furthermore, we analyzed how user activity levels affect prediction uncertainty in our online system, as shown in Table 4.
Table 4. Prediction performance across user activity levels in online environment
User Activity Level IQR UAUC MAE
Low Activity 0.289 0.620 0.200
Medium Activity 0.274 0.634 0.195
High Activity 0.273 0.643 0.193
These findings align with our theoretical understanding: (1) less active users exhibit higher prediction uncertainty (larger IQR); (2) prediction accuracy improves as user activity increases, likely due to richer historical data; and (3) even highly active users show inherent prediction uncertainty, underscoring the universal need for our distribution-aware approach regardless of user engagement level.
This empirical evidence from online data strongly supports our quantile-based framework and provides additional justification for the three inference strategies introduced in Section 3, each designed to handle different aspects of prediction uncertainty.
@@ 4.2. Distribution Characteristics Analysis
Figure 3. The diversity performances of CQEDQC compared with baseline online.
4.3. Offline Experiments (RQ2 and RQ3)
Table 5. Comparison of CQE with other methods across different backbones in the user interest prediction task. Additional results for the DCNV2 backbone are provided in Appendix A. Boldface highlights the best-performing method for each backbone.
Backbone Method KuaiRand WeChat
GAUC nDCG@1 nDCG@3 nDCG@5 GAUC nDCG@1 nDCG@3 nDCG@5
DeepFM PCR + MSE 0.592 0.411 0.465 0.502 0.537 0.565 0.572 0.578
PCR + CE 0.635 0.435 0.486 0.521 0.553 0.584 0.586 0.591
PCR + CQECDE(Ours) 0.638 0.438 0.488 0.523 0.593 0.654 0.649 0.649
WTG + MSE 0.601 0.417 0.470 0.506 0.537 0.578 0.579 0.585
WTG + CE 0.647 0.440 0.491 0.526 0.545 0.565 0.569 0.576
WTG + CQECDE(Ours) 0.654 0.444 0.495 0.529 0.577 0.621 0.622 0.625
$\text{D}^{2}$o + MSE 0.606 0.416 0.473 0.509 0.528 0.578 0.577 0.580
$\text{D}^{2}$o + CE 0.656 0.447 0.496 0.530 0.531 0.570 0.571 0.576
$\text{D}^{2}$o + CQECDE(Ours) 0.662 0.452 0.500 0.533 0.557 0.635 0.626 0.625
AutoInt PCR + MSE 0.595 0.414 0.467 0.503 0.536 0.565 0.572 0.578
PCR + CE 0.639 0.437 0.488 0.522 0.554 0.579 0.583 0.588
PCR + CQECDE(Ours) 0.640 0.441 0.489 0.524 0.593 0.655 0.649 0.649
WTG + MSE 0.602 0.419 0.470 0.507 0.536 0.568 0.573 0.578
WTG + CE 0.648 0.441 0.492 0.526 0.546 0.563 0.570 0.576
WTG + CQECDE(Ours) 0.652 0.448 0.495 0.528 0.576 0.631 0.625 0.627
$\text{D}^{2}$o + MSE 0.610 0.420 0.474 0.510 0.527 0.582 0.580 0.583
$\text{D}^{2}$o + CE 0.658 0.447 0.497 0.531 0.531 0.570 0.571 0.576
$\text{D}^{2}$o + CQECDE(Ours) 0.661 0.455 0.500 0.534 0.557 0.635 0.626 0.624
While our online A/B tests demonstrated the practical impact of CQE in real-world scenarios, offline experiments allow us to conduct more controlled and detailed analyses of our approach. Our offline experiments focus on two closely related tasks: watch time prediction and user interest prediction. Together, these tasks provide a comprehensive perspective for evaluating the effectiveness of the CQE framework in recommendation systems.
Watch time prediction directly captures the duration of user engagement with content, a key indicator of user involvement. However, predicting watch time alone may not fully capture user interest. Therefore, we introduce the user interest prediction task, accounting for video duration bias to offer a more refined measure of user interest. These two tasks complement each other: watch time prediction offers direct behavioral forecasting, while user interest prediction helps us understand the motivations behind these behaviors.
Following their publicly available code, we adopted the same network architecture and training configuration as TPM (Lin et al., 2023) for watch time prediction and $\text{D}^{2}$o (Zhao et al., 2023) for user interest prediction.
4.3.1. Watch-Time Prediction.
In this task, our primary objective is to accurately predict the duration of user watch time.
Datasets. Following TPM (Lin et al., 2023), we used two public datasets: Kuaishou (collected from Kuaishou App [^1] https://kuairec.com/) and CIKM16 (from CIKM16 Cup  [^2] https://competitions.codalab.org/competitions/11161) for our experiments. While CIKM16 is primarily an e-commerce search query dataset, we include it to demonstrate the potential generalizability of our CQE approach across different recommendation contexts. The prediction of page dwell time in e-commerce shares similarities with video watch time prediction in terms of modeling user engagement duration, although we acknowledge the differences in content type and user behavior patterns. In CIKM16 dataset, each item in the session is used as a single feature for input. Kuaishou dataset contains 7,176 users, 10,728 items, and 12,530,806 impressions; and CIKM16 dataset contains 310,302 sessions, and 122,991 items and the average length of each session is 3.981.
  • MAE: This metric is a typical measurement to assess the accuracy of the regression. Denote the prediction as $\hat{y}$ and the true watch time as $y$,
    (9) $\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{y}_{i}-y\|.$
  • XAUC: It evaluates if the predictions of two samples are in the same order as their true watch time. It aligns well with the ranking nature of recommendation systems. In practice, the relative order of predictions often matters more than their absolute values, making XAUC particularly relevant to our research goals.
Baselines. Six state-of-the-art methods for watch time prediction were selected for comparison, including WLR (Covington et al., 2016), D2Q (Zhan et al., 2022), OR (Niu et al., 2016), TPM (Lin et al., 2023), DML (Zhang et al., 2023) and CREAD (Sun et al., 2024). The results of the first four methods are sourced from the TPM paper, while the latter two are based on our own reimplementations, as their official code is not publicly available.
4.3.2. User Interest Prediction.
This task accounts for video duration bias to extract user interest from watch time. By doing so, we consider not just how long a user spent watching a video, but also how this time relates to the total video length, thus more accurately reflecting the user’s true level of interest.
Following $\text{D}^{2}$o (Zhao et al., 2023), in detail, we define the user interest for a given user-video pair $(u,v)$ as:
where $d$ is video duration and $w$ is watch-time. We adopted the same training configuration as $\text{D}^{2}$o and used the classical deep recommendation model DeepFM (Guo et al., 2017), the state-of-the-art recommendation model AutoInt (Song et al., 2019) and DCNV2(Wang et al., 2021) as our backbone recommendation model.
where $\sigma$ is the sigmoid function and $r$ is the user’s interest defined by PCR(Play Completion Rate), WTG(Watch Time Gain) (Zheng et al., 2022) or $\text{D}^{2}$o(Debiased and Denoised watch time Correction) (Zhao et al., 2023). Following $\text{D}^{2}$o, in PCR and WTG, we treat all samples with less than 5 seconds of watch time as 0 values after calculating the value of labels. This can help remove the noise in watch time.
By default, we set the number of quantiles $N$ to 100. The value of $\tau_{low}$ is selected empirically from 0.2, 0.25, and 0.3. Similarly, the value of $\tau_{high}$ is selected empirically from 0.5, 0.6, 0.7, and 0.8.
Table 6. Comparison between CQE and other approaches in watch-time prediction task. Boldface means the best-performed methods.
Methods Kuaishou CIKM16
MAE XAUC MAE XAUC
WLR 6.047 0.525 0.998 0.672
D2Q 5.426 0.565 0.899 0.661
OR 5.321 0.558 0.918 0.664
TPM 4.741 0.599 0.884 0.676
DML 4.455 0.607 0.829 0.693
CREAD 4.458 0.608 0.831 0.691
CQECDE(Ours) 4.437 0.610 0.823 0.694
4.3.3. Experimental Results
We summarize the results as follows:
Comparison between CQECDE and other methods (RQ2): We compared the performances of different approaches in the watch-time prediction task, and the results are listed in Table 6. Our approach demonstrates superior performance on both metrics compared to other mtehods, thereby emphasizing the advantages of employing quantile modeling technique. Additionally, the consistent behavior between MAE and XAUC metrics also verifies the feasibility of the watch-time estimation serving as ranking metrics. As for the user interest prediction task, we compare different frameworks across backbone models (DeepFM, AutoInt and DCNV2) and various label designs (PCR, WTG, and $\text{D}^{2}$o) and present the results in Table 5 and Appendix A. Our proposed CQECDE consistently outmatches the alternatives in all cases, indicating the robustness and effectiveness of CQECDE. In terms of the optimization framework, CE generally performs better than MSE, indicating the correctness of including the ordinal categorical information as guidance. And CQECDE can improve over CE across all the designs of user interest metrics (PCR, WTG, and $\text{D}^{2}$o), which means that the proposed framework is generalizable to different label settings.
The Effect of Hyper-Parameters in CQECDE (RQ3): To better investigate the characteristics of the proposed CQE framework, we further conduct an ablation study on the number of quantiles $N$ by varying its value from 1 to 500. Theoretically, larger $N$ generates a more accurate approximation of the truth expectation and in turn achieves better recommendation performance in general. This is attributed to the fact that more quantiles yield a distribution that closely mirrors the actual one. As shown in Figure 4, model performance in the watch-time prediction task improves with an increasing number of predicted quantiles. For user interest prediction (Figure 5), performance remains relatively weak when the number of quantiles is below 10. Beyond this threshold, results fluctuate around 0.663, indicating that unlike watch-time prediction, a higher number of quantiles does not necessarily enhance performance. This discrepancy suggests a misalignment between the training objective and the user interest labels in the test set. Nevertheless, increasing $N$ generally improves prediction accuracy under the Conditional Expectation strategy.
In summary, our offline experiments comprehensively demonstrate the superiority of the CQE method in predicting user behavior and interest through these two complementary tasks. The watch time prediction task validates CQE’s accuracy in direct behavioral prediction, while the user interest prediction task further proves CQE’s ability to effectively capture more complex user preferences. The combination of these tasks not only validates the effectiveness of our method but also highlights the flexibility and adaptability of the CQE framework in addressing different yet related challenges in recommendation systems.
Figure 4. The performances of CQECDE with various number of quantiles in CIKM16 Dataset for predicting watch time.
[…]
A woman and a girl in white dresses sit in an open car.
Figure 5. The performances of CQECDE with various number of quantiles in KuaiRand Dataset for predicting user interest.